OpenStudy (trojanpoem):

Z is a complex number, a is a constant. Find the maximum and minimum value of the complex number Z notice that :

OpenStudy (trojanpoem):

\[\left| \frac{ 1 }{ Z } + Z \right| = a\]

OpenStudy (trojanpoem):

a is not equal to zero

OpenStudy (trojanpoem):

It's a complex number , you can assume it's A + Bi

OpenStudy (trojanpoem):

It is.

OpenStudy (trojanpoem):

No problem :D

OpenStudy (phi):

the the max and min of what ?

OpenStudy (trojanpoem):

of the complex number Z

OpenStudy (phi):

how do we define the max of a complex number? do you mean the max magnitude ?

OpenStudy (trojanpoem):

You can the maximum of A , B ( A+ Bi ) is the complex number.

OpenStudy (trojanpoem):

Yeah, the magnitude

OpenStudy (phi):

you could write Z= a exp( i x) so Z+1/Z = a exp(ix) + (1/a) exp(-ix) and the magnitude squared is | Z+1/Z |^2 = (a exp(ix) + (1/a) exp(-ix))(a exp(-ix) + (1/a) exp(ix)) which might lead to a solution

OpenStudy (trojanpoem):

Hmm, do you mean that e^x ? Never heard of it :(

OpenStudy (phi):

no. I meant a complex number in polar coords is magnitude * exp( i* theta) btw, I should have used r instead of a, to not confuse the magnitude with the "a" given in your problem.

OpenStudy (phi):

exp(i theta) means \( e^{i \theta}\)

OpenStudy (trojanpoem):

Oh, Got it

OpenStudy (trojanpoem):

Ok now we ave e^thetai + e^-thetai